OpenStudy (anonymous):

solve x(x+5)=14 a.-7,2 b.-5,4 c.-3,6 d.-2,7

OpenStudy (ash2326):

(x(x+5)=14)(x x^2+5x-14=0 factor 14 such that the difference or sum of them is 5 that is 7 and 2 5 =7-2 x^2+7x-2x-14=0 x(x+7)-2(x+7)=0 (x-2)(x+7)=0 x=2 and -7 option a

hero (hero):

x^2+5x - 14 = 0 x^2+7x - 2x - 14 =0 x(x+7)-2(x+7) = 0 (x+7)(x-2) = 0 x = -7,2